OpenStudy (anonymous):

f(d)= (d^4-d+1)1/3 derivative function, please help to answer this. show me the proper wat to solve to easy to understand!

OpenStudy (perl):

interesting problem :)

OpenStudy (anonymous):

@perl can you help sir?

OpenStudy (anonymous):

@ganeshie8 sir can you help me?

OpenStudy (perl):

if i understood the problem i could help :)

OpenStudy (anonymous):

please help me. :((

OpenStudy (anonymous):

You are given the derivative, or you want to find the derivative?

OpenStudy (anonymous):

find sir

OpenStudy (anonymous):

@wio sir please help me

OpenStudy (anonymous):

This?\[ f(d)= (d^4-d+1)^{1/3} \]

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Do you know the chain rule?

OpenStudy (anonymous):

no sir

OpenStudy (anonymous):

can you show me the proper way to solve?

OpenStudy (anonymous):

Hmm, to be honest, I'm not sure how to do this without using the chain rule.

OpenStudy (anonymous):

Are you sure they did not teach the chian rule?

OpenStudy (anonymous):

no my prof teach how to solve thabuy im absent always because of my nightshiht callcenter parttime

OpenStudy (anonymous):

so i dont know how to solve that? :'(

OpenStudy (anonymous):

@strawberryswing help

OpenStudy (anonymous):

@sammixboo help

OpenStudy (anonymous):

@Zale101 help

OpenStudy (anonymous):

Okay, what rules do you know?

OpenStudy (anonymous):

Ahhhhhh math

OpenStudy (anonymous):

Run!

OpenStudy (anonymous):



OpenStudy (anonymous):

 @wio

OpenStudy (anonymous):

Okay, now consider that: \[ f(d) \times f(d)\times f(d) =d^4-d+1 \]

OpenStudy (anonymous):

You can differentiate both sides, using the product rule and power rules.

OpenStudy (anonymous):

So \[ 3(f'(d)\times f(d)\times f(d)) = 4d^3-1 \]Does this make sense?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Now we plug in \(f(d) = (d^4-d+1)^{1/3}\)

OpenStudy (anonymous):

So \[ f'(d) \cdot 3(d^4-d+1)^{2/3} = 4d^3-1 \]

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

so sir what is the final answr

OpenStudy (anonymous):

Can you solve for \(f'(d)\)? That is what we originally wanted.

OpenStudy (anonymous):

i dont knw how to solve that sir

OpenStudy (anonymous):

sir?

OpenStudy (anonymous):

It's algebra... divide the part that isn't \(f'(d)\).

OpenStudy (anonymous):

@wio 0?